confused the answer i got was -300 but confused on the second part.For the following equation, state the value of the discriminant and then describe the nature of the solutions.-5x^2+10x-20=0 What is the value of the discriminant?
Answer by stanbon(75887)

For the following equation, state the value of the discriminant and then describe the nature of the solutions.-5x^2+10x-20=0 What is the value of the discriminant?
------------
b^2-4ac = 10^2-4*-5*-20 = -300
----
Since the discriminant is negative,
there are 2 unequal complex solutions.
===========================================
Cheers,
Stan H.